### v1

# Cultural Transmission, Technology, and Treatment of the Elderly

Abstract
We discuss the interrelationship between the treatment of the elderly, the nature of production, and the transmission of culture. Respect for the elderly is endogenous. Parents cultivate an interest in consuming culture in their children; when they are older, children compensate their elders proportional to the degree to which their interests were previously cultivated. We show that this model is functionally equivalent to one in which cultural goods are transferred across generations. We focus on the relative well-being of the elderly and use the model to explain patterns in their relative well-being across societies. An important theme is that the cultivation of culture and norms for the respect and support of the elderly bear a nonlinear relationship with many economic variables, such as capital and or land intensity in production. We also discuss the interaction of property rights with production, assets such as productive resources, and relative treatment of the elderly. Insecurity of some types of property rights, such as rights over output, may benefit the elderly, while secure rights over productive resources may also benefit the elderly. We discuss how the elderly could be affected by demographic, technological and policy changes in both developing and developed economies.
1 Introduction
Aging and the elderly have always been with us. Although this has always and everywhere been true, [^1] A finding of recent anthropological research is that living into old age is common among hunter-gatherers, and was likely common even among early Homo Sapiens (Gurven and Kaplan, 2009). In fact, long lives have been hypothesized to be an evolutionary outcome of the hunter-gatherer lifestyle (Kaplan and Robson, 2002, 2003) there are pronounced differences in the way aging is experienced around the world. In modern and highly developed economies, the elderly are typically able to provide for their own well-being and also receive government-provided benefits. [^2] A United Nations report suggests that, in developed countries, public transfers are the most common means of elderly support in Europe and South America, while in Southeast Asia, the United Kingdom, and the United States, elderly support is mainly financed by asset income. (United Nations Department of Economic and Social Affairs, 2019) In less developed and more traditional societies, including preliterate societies, there is great diversity in the way elderly are cared for and variety in the societal roles the elderly play. The elderly may receive special treatment for the particular knowledge they possess or also for their role as keepers of history and cultural stores. The elderly may benefit from specialized property rights or from the development of property rights in general, as is the case when the elderly gain ownership over productive assets such as land. Sometimes, the elderly may simply take advantage of having lived longer than others, allowing for the accumulation of productive assets and experience.
Demographic and technological changes exert ongoing pressure on the institutions and cultural practices that shape the aging experience. People now live longer and have fewer children than ever before, increasing both the relative and absolute size of the elderly population virtually everywhere. In 1950, there were three people under the age of 5 for every person over the age of 65; projections suggest that by 2050 the opposite will be true (World Health Organization, 2011). Worldwide changes in work, consumption, and savings patterns have consequences for everyone, but these changes have special meaning for the elderly. Some popular writing, such as Diamond (2012, Chapter 6), has drawn attention to aspects of the changing situation of the elderly, and policy makers have begun to craft legislation designed to address the changing situation of the elderly in modern society in some parts of the world. [^3] See https://theweek.com/articles/462230/how-elderly-are-treated-around-world.
Across the development spectrum, and particularly in less developed economies, the elderly play a critical role in the maintenance and transmission of culture. Therefore, it is natural to ask the following. What implications might the changes that profoundly impact the elderly have for cultural transmission and retention? [^4] As one report on global aging puts it, “…some countries may grow old before they grow rich.” (World Health Organization, 2011, p. 5) Much of the growth of the elderly population in the coming decades is expected to occur in less developed countries, where the condition of the elderly is closely linked with tradition and culture. [^5] World Health Organization (2011) estimates that the world’s population 65 or older will grow from 524 million in 2010 to 1.5 billion in 2050, most of this growth occurring in developing countries. United Nations Department of Economic and Social Affairs (2019) estimates that by 2050, 1 in 6 people in the world will be over the age of 65.
We construct a simple model that provides insight into the forces driving the diversity of elderly treatment across societies and across the spectrum of development, from hunter-gatherer societies to modern economies. Since the model sheds light on some of the aspects of material life that have led to cultural practices that support the elderly, it provides insight into how technological and demographic change impact the way elderly are treated. At the same time, the model provides insight into how elderly treatment and cultural transmission are intertwined and therefore offers an understanding of how transmission of culture itself interacts with technological and demographic change.
2 Literature
Formal analysis of the condition of the elderly has attracted the attention of economists at least since mathematical models of growth became an integral part of economics. Aging, savings behavior, and capital accumulation are intimately connected in many, if not most, of the core dynamic economic models. The life cycle model of Modigliani (1966) is explicitly built on the relationship between aging, savings behavior, and accumulation. The overlapping generations (OLG) model developed in Samuelson (1958) and Diamond (1965) also intimately binds savings, demographics, and accumulation.
The OLG model has proven to be useful in analyzing the interrelationship between capital accumulation, aging, and family structure, and has also been used to study the role that inter vivos transfers play in an economy, forging a direct link between generations. [^6] A review of literature on altruism and inter-generational transfers is Michel et al. (2006). Boldrin and Jones (2002) and Nishimura and Zhang (1995) reverse the usual parent-child altruism of Barro and Becker (1989), and instead study situations in which children are altruistic towards their parents. These models inform elderly treatment; in an environment with child-to-parent altruism, transfers to the elderly emerge naturally, as do investments in fertility and children’s education, which may be altruistic but may also result from self-interested behavior on the part of the elderly, as such investments may enhance subsequent transfers to the elderly.
Some research, following models of strategic interfamilial behavior like Bernheim et al. (1985), posits that intergenerational transfers to and from the elderly may play a critical strategic role in the maintenance of the family fabric. Cigno (1993) show how transfers from the current working-age population to both young and old may be part of a self-enforcing set of family rules; a chief finding of Cigno (1993) is that the existence of such rules is sensitive to the existence of capital markets and social support systems. [^7] Kaplan and Robson (2003) have suggested that an important aspect of human evolution might require such intergenerational capital markets. As Laferrére (1999) documents in her survey of models of intergenerational transmission, there are a wide variety of models in which strategic motives both in transferring wealth between family members of different generations and actively encouraging altruistic behavior between family members emerge. A prominent example is Morand (1999), who discusses the relationship between fertility and growth in a model in which the elderly make fertility decisions with an eye on the support of the elderly, linking the literature on demographic transition (Galor and Weil, 2000) and the support of the elderly. Blackburn and Cipriani (2006) develop a model in which societies transition from economies characterized by high fertility rates, low levels of development, and intergenerational transfers that flow from young to old, to economies with low fertility rates and high levels of development, with transfers flowing from old to young.
A literature focusing on cultural transmission of familial norms through parental socialization efforts and the presence of coincident traits in the population, following the approach introduced in Bisin and Verdier (2000, 2001), has also been applied to shed light on the evolution of elderly treatment. Olivera (2013) studies a model in which agents have a genetic disposition to care for the elderly (or not), and shows how support for old age relates to fertility decisions. In this type of model, norms of elderly support may persist, even as fertility declines and life expectancy increases. Varvarigos (2021) explores a model similar in spirit to Olivera (2013) in which adults influence children’s preferences to be like their own, but with a focus on the implications for development. Varvarigos’s model produces equilibria in which low levels of economic development, strong family ties, and intergenerational transfers coexist; the model is also capable of producing equilibria in which intergenerational transfers are not important, family bonds are not as strong, and the economy functions at higher levels of development. Cigno et al. (2017) and Cigno et al. (2020) combine marital classification and trait transmission with aspects of intergenerational self-enforcing contracts originally developed in Cigno (1993). This work shows how changes in education levels and wages can reduce the prevalence of family contracts focusing on the elderly.
The Bisin and Verdier (2000, 2001) model style has also been used to study the relationship between social support systems and elderly treatment. Cremer et al. (2013), for example, presents a model in which parents spend time with children as a means to gain support for a possible future disability and study the interaction of support with policy. Ponthiere (2013) discusses a replacement effect in which state-provided long-term care can modify the transmission of values and affect family altruism.
Although our model is similar in spirit to these models, the basic mechanism underlying culture transmission and how it relates to support for the elderly is different. In our model, middle-aged agents work to cultivate an interest in their children in the consumption of a cultural good provided by the elderly, as is reminiscent of Cozzi (1998). Adopting this approach firmly focuses our paper on the role in which cultural transmission is related to other savings decisions that the agents might make, rather than working through a secondary channel in which agents care about the propagation of their genetic traits. It also forms a direct link between the appreciation and maintenance of culture and the treatment of the elderly. Under some circumstances, our approach is then isomorphic to an approach in which agents invest in cultivating an interest in culture as a means of storing wealth for consumption in the future. Since current middle-aged will be elderly in the future, if they cultivate interest in cultural consumption among their children in the present, future consumers of their cultural goods may in part sustain them in the future.
There has been extensive empirical work on altruism toward the elderly. Early work on intergenerational transfers within the family found that forming hard conclusions about the motivations driving transfers is elusive (Laferrére, 1999), as purely altruistic motives for giving are difficult to separate from self-interested motives. In a recent example, Klimaviciute et al. (2017) apply data from a large European dataset on children’s and parents’ behavior concerning intergenerational transfers to test models of altruism towards the elderly, to determine how support policies could interact with the interfamilial provision of elderly care. Their findings support “moderate altruism” as the driving force behind inter vivos transfers to the elderly.
Therefore, there have been a variety of different approaches to studying the evolving nature of elderly treatment and how elderly treatment interacts with strategic, altruistic and technological aspects of the economy. Our approach extends these approaches in a way that melds strategic and altruistic aspects of elderly treatment, and that also explicitly links economic growth, the development of property rights, and the nature of production to the treatment of the elderly. Most previous research [^8] But not all. Varvarigos (2021), for example, is a general equilibrium analysis. relies on a partial equilibrium approach and does not seek to explore deeper trends in elderly treatment, production technology and economic growth in a way that incorporates a typical growth model such as that espoused in Acemoglu (2009, Chapter 9).
Furthermore, the existing literature does not explicitly focus the analysis on the treatment of the elderly. The cross-cultural ethnographic literature points out several critical factors of the economy driving the treatment of the elderly, such as the origins and development of property rights. Our approach allows us to build a model capable of replicating this evidence, while along the way developing working measures of elusive concepts such as “elderly treatment” and “elderly prestige.” Finally, our model tightly links technical change and elderly treatment and allows us to see how preferences, elderly treatment, and technology co-evolve.
3 The Ethnology of Elderly Treatment
3.1 Cross-cultural research
The logical place to begin a discussion of cross-societal analysis of the treatment of the elderly is the seminal work of Simmons (1970) (originally written in 1945), who surveyed the treatment of the elderly in a sample of 71 different societies. Simmons’ sample included hunter-gatherers, pastoral and nomadic herding peoples, and subsistence agriculturalists from all over the world. In constructing these data, Simmons aimed to create a sample of societies that were at once technologically, geographically, and environmentally diverse. [^9] Simmons’s approach mirrors that employed by G. P. Murdock and his co-researchers in the construction of the Ethnographic Atlas (Murdock, 1969), and Murdock and White in their construction of the Standard Cross-Cultural Sample (Murdock and White, 1969), and also resembles the data used in Pryor (1973). See Baker and Jacobsen (2007b), Baker and Jacobsen (2007a), or Baker and Miceli (2005) for an application and further description of cross-cultural data within the context of economic modeling. This sample allowed Simmons to develop a fairly comprehensive picture of regularities in elderly treatment across societies and thus form a picture of how the role of the elderly changes along with the nature of subsistence and societal complexity. In addition to setting the stage for subsequent work on cross-cultural treatment of the elderly, Simmons’s survey also provides a useful snapshot of elderly treatment around the world before the onset of global modernization. The following subsections organize and explain some of the key results of this exercise. Our summary of this information and the organization of our summary also relies on the review of Simmons’ work in Holmes (1976) and the discussion of Arnhoff et al. (1964).
3.1.1 Food Sharing and Taboos
In societies without well-developed property rights (perhaps subsistence hunter-gatherer societies), practices such as communal food sharing and food taboos provide consumption security for the elderly. Elderly consumption becomes more secure as property rights in land and in other property, such as livestock, develop.
Most cultures earning subsistence by hunting and gathering also usually have some system for food sharing. The existence of sharing has been attributed to a wide variety of motives - insurance, exchange, even resource management - making it difficult to surmise “the” reason output is shared. [^10] See Baker and Swope (2021) for an overview of the many theories of sharing among hunter-gatherers. Kelly (2013, Chapter 6) has an extensive discussion of sharing across hunter-gatherer populations. Nonetheless, a common feature of food sharing is that the elderly tend to be included, so food sharing provides food for the elderly. Thus, to use the terminology of Lee (2003), sharing redistributes “vertically” from adults to both the elderly and children, as well as “horizontally” - that is, between hunters. It is also true that many hunter-gatherers share access to resources, so it seems that the lack of individual ownership rights over the food output that is produced is correlated with the lack of ownership over land and other resources. [^11] This of course does not mean hunter-gatherers aren’t territorial at the group level; see (Baker, 2003).
The practice of tabooing food in a way advantageous to the elderly, referred to as “Victual Deference” by Silverman and Maxwell (1978), is common among hunter-gatherers living in a wide variety of environments. Simmons describes taboos as “…a social expedient of particular advantage to the aged…” and also notes that “…while it cannot be demonstrated that these oldsters invented such social expedients in behalf of their own palates, many of the tabooed foods were regarded as quite tasty, and frequently the aged have been observed manipulating such taboos to their personal advantage.”(Simmons, 1970, p. 26) Food taboos are a useful illustration of a core principle embodied in the last item on the list, as they are very likely ingrained in individuals through conscious effort. They therefore represent a telling example of a cultural practice that is taught and transmitted and, at the same time, may actually serve no other practical purpose than to secure subsistence for the elderly. Simmons gives food tabooing examples from a diverse group of societies around the world, including Native American groups such as the Greenland Inuit, Omaha, and Pomo; the Iban of Borneo; and the Arunta from the Australian desert, to name a few.
Examples beyond Simmons’s work of taboos are easy to find. Food taboos feature prominently among the Ju/’hoansi, a hunting and gathering people that inhabit the Kalahari desert in Namibia and Botswana, who have been the subject of a wealth of research. [^12] In English, this would be something like ‘juntwasi.’(Rosenberg, 2009) In earlier writings, they were known as the !Kung. See, e.g., Lee (1979). For example, “elders are felt to have special powers that permit them to eat certain foods (e.g., ostrich eggs) considered too dangerous for younger people to consume.” (Rosenberg, 2009, p. 33). McArdle and Yeracaris (1981) provide another example from the Chukchee people of Siberia that they attribute to Bogoras (1904), in which it is believed that Reindeer milk is harmful to young people. An interesting variation on this theme comes from the Pukapukan people of the Cook Islands, where there exists a “culturally conditioned system of both food sharing and food taboos” (Simmons, 1970, p. 10) that favored the elderly, in which the elderly controlled the distribution of food and consequently were able to direct to themselves some of the most nutritious animal parts, such as belly fats and internal organs.
3.1.2 Property rights
While in the case of food sharing, it might be said that ill-defined property rights benefit the elderly, Simmons observes that “property rights have been lifesavers for the aged,” (Simmons, 1970, p. 36), implying that the definition of other types of property rights benefit the elderly. Simmons overlaps a few different concepts of property rights, and the distinction is important. As discussed in the previous section, nonexclusive rights over food help the elderly, but it is also the case that ownership over assets, as it emerges, may benefit the elderly.
Simmons finds that among people who rely on agriculture or animal husbandry (or both) for subsistence, the elderly are also afforded respect as among hunter-gatherers, but for seemingly different reasons, as in these cases respect likely owes to accumulation of wealth. Simmons documents the accumulation of assets such as working animals and livestock among the elderly in societies such as Navajo (Southwestern North American Desert), Chukchi (Siberia near the Bering Strait), Yakut (Russian far east), and Akamba (Kenya), to name a few. Peoples who rely primarily on agriculture coupled with land ownership also seem to feature respect for the elderly coincident with accumulation of assets. As (Simmons, 1970, p. 44) indicates, “Firmly entrenched property rights which enhanced the powers of the aged have also been very common among all cultivators of the soil.”
3.1.3 Knowledge and the Elderly
The elderly often provide valuable services securing them a livelihood, and often provide administrative services, which can convey power and prestige to the elderly. Among the less developed peoples in Simmons’ sample, the elderly may receive payment for the provision of services that confer something of material value, such as childcare or food processing, or even technical knowledge or administrative know-how. The status of the elderly increases and treatment improves with the importance of tradition in society. Elderly status also improves with the amount of specialized knowledge the elderly have.
But Simmons also discusses how the elderly also often provide purely cultural goods and services in exchange for compensation. For example, elderly Navajo were compensated for their knowledge of magic songs and charms, while among the Pomo and the Lengua, elderly medicine men earned fees for performing important rites (Simmons, 1970, p. 37-38). Thus, some types of property and associated property rights appear to be established by cultural conditioning, as is the case with rights over songs and charms, while at the same time, other types of cultural conditioning alter or expand property rights in ways that aid the elderly. As an example, rules for sharing game or food tabooing might be thought of in this way.
The belief that the elderly possess magical powers is common among subsistence hunters and hunters, instilling fear and awe of the elderly. The elderly are often believed to be endowed with the power to bring and cure disease, and these beliefs are actively cultivated. The cultures in which such beliefs were prevalent operate in widely varied material circumstances; the Yahgan (Tierra del Fuego), the Arunta (Australia), the Greenland Inuit, the Pomo (California Desert), and the Witoto (Northern Peru). Rosenberg (2009, p. 34) notes that the Ju/’hoansi hold the belief that some elderly have powers of healing, which are due to their experience allowing them to use certain mystical forces without being overwhelmed by them.
3.1.4 Inculcation
Attitudes about the elderly and cultural practices directed towards the elderly are consciously nurtured in the young, sometimes at considerable cost to parents. Simmons (1962) is clear to indicate that these forms of elderly treatment arise due to the application of costly resources and effort. Simmons notes that “Respect for old age has resulted from social discipline,” (Simmons, 1970, p. 50), and, writing more expansively (his italics):
“…while human beings can be as indifferent to the dependency needs of their ancestors…they can also learn, be taught, inspired, or impelled to respect, succor, and sustain their elders…these patterns of group and filial responsibility are ‘passed on’ from generation to generation…This learned behavior is called culture.” (Simmons, 1962, p. 37)
Thus, some degree of cultural conditioning allows the elderly to consume and control resources that they would not be able to otherwise.
3.2 Elderly Status and Cultural Complexity
Similar conclusions also emerge in subsequent cross-cultural research on the treatment of the elderly. Press and McKool Jr. (1972, p. 298) give a brief summary of the “ostensibly universal generalizations about status of the aged” offered by Cowgill and Holmes (1971), which, more or less, follow those of Simmons (1970). These generalizations include that the status of the elderly is high when knowledge accumulation is important in the economy, when the elderly can hold property rights, and when the elderly can perform economically meaningful tasks. In fact, Press and McKool Jr. (1972, p. 299), in developing their own list of determinants of elderly status, came to the conclusion that (their italics) “…our independently derived determinants coincided with Cowgill’s.”
Our discussion has so far focused on the elderly among hunter-gatherers and cultures that rely on subsistence agriculture or animal husbandry for subsistence, but much subsequent cross-societal research has focused on how elderly treatment evolves as societies progress further along some dimension of technological complexity, which seems to encompass dimensions such as the definition of property rights, social complexity, and technological change in the basic mode of production of the economy. In this research, elderly treatment becomes enmeshed in research on the complexity of familial structure, but a broad trend in elderly treatment that has emerged from this line of inquiry is that it is difficult to define an exact trajectory of elderly treatment as the economy becomes more complex. Put differently, elderly treatment varies in a non-linear fashion with the level of development of society. One difficulty plaguing this branch of research is the reliance on overly vague terminology, where terms such as “status” and “good treatment” are used without formal definition.131313 Dowd (1981) and Balkwell and Balswick (1981) note the problems caused by the poor definition of relevant terms. What does it mean to say, for example, that the elderly are treated well in a subsistence hunting and gathering community but not all that well treated in a modern industrialized economy? Some researchers may claim that the status of the elderly is increasing in the level of societal development, and while other researchers may disagree with this, the disagreement may really be about how status is defined and what the relevant range of development is.
Despite some confusion about these terms, an important characterization is due to the so-called “curvilinear hypothesis,”(Blumberg and Winch, 1972) as originally developed in Nimkoff and Middleton (1960).141414(Lee, 1996) is a good assessment and an overview of the hypothesis. See also Lee and Kezis (1981) and Baker and Miceli (2005). According to the hypothesis, there is a nonlinear relationship between social complexity and the extent of the family. Those societies at the lowest levels of development (that is, hunter-gatherer societies) and those at higher levels of development (intensive agriculturalists and industrial societies) tend to have smaller families. Those societies “in the middle” tend to have the most expansive family structures, and, as Lee (1979) documents, the elderly tend to have a high status in these familial structures. This literature considers the nexus of things documenting the importance of familial structure in elderly treatment, but often confounds what are really two different things: the elderly being treated preferentially because of a cultural predilection to do so, and the elderly commanding respect because they have accumulated resources in the past. This is at the crux of the matter: How do accumulation of assets and cultural transmission relate to each other?
3.3 Modernization and the Elderly
As globalization and technological change continue, as population growth slows, and as the global population ages, it is important to understand how technical change interacts with cultural practices relevant to the elderly. Although there seems to be some reason to believe that industrialization is detrimental to the well-being of the elderly (Dowd, 1981, e.g.), the relationship between different aspects of industrialization and the well-being of the elderly has turned out to be complex. Understanding these relationships is even important in modern societies, and the study of such problems naturally leads one to wonder how the institutions and cultural practices that provide care for the elderly are impacted by modern institutions, such as a social security system. These questions become more important when considering that much of the world’s population will be transitioning from lower to higher levels of development in the coming decades. Koda and Uruyos (2018) for example, find evidence that public transfers crowd out private care for the elderly and also that the development process, which leads to lower fertility and smaller families, tends to reduce family support for the elderly, offering some support for the theories presented in Koda and Uruyos (2018) and Cigno (1993). In contrast, Ko and Möring (2021) find in their study of pensions in rural China that receiving a public pension accompanies an increase in both well-being and the degree of intergenerational exchange. Some research (Kodate and Timonen, 2017) has described some of the complexity of the interaction between public funding, paid health care workers, culture, and family care.
Therefore, elderly treatment and status interact with the standard desire to smooth consumption in rather complex ways. In general, the desire to push consumption of current resources into the future may result in improving treatment through cultural means, accumulation of assets over time, or both. How exactly this happens depends on the nature of production and other basic characteristics of the economy.
Developing a model in which these trade-offs can be explored turns out to shed much light on these issues and maybe even lends insight into how elderly treatment will evolve in the coming decades. A model should allow us to better understand the interplay between elderly treatment and the underlying attributes of the economy and also generate predictions that capture some of the complexity in the ethnographic record. In the next section, we present such a model. An additional useful aspect of our model is that it requires defining some of the often used concepts in the empirical literature that remain vague, such as ‘status’ or ‘prestige’, or even what it means for the elderly to be ‘treated well.’
4 Theoretical Framework
The model is a standard infinite-horizon overlapping generations model, where at any given time young, middle-aged, and elderly agents co-exist. Agents make decisions solely in middle age. When young, agents do not earn any utility but can be inculcated to care about culture. In our base model, current middle-ageds cultivate a taste for giving gifts to the elderly in the young. When young people reach middle age, they are willing to transfer resources to the elderly, where their willingness to transfer resources and earn ‘warm glow utility’ as in Andreoni (1989) depends on the degree to which appreciation was cultivated in agents currently in middle age when they were young. [^15] In this respect, our model resembles one in which cultural heritage and society’s devotion to it is endogenous, in a similar fashion to Cozzi (1998). Cozzi’s model, however, is built on the premise that beliefs about the value of culture are self-fulfilling and are not actively cultivated.
The resulting model of endogenous transfers to the elderly is observationally equivalent to a model in which appreciation for a cultural good is cultivated in the young and the cultural good is supplied by the elderly. Thus, the endogenous transfer, warm-glow formulation of the model captures an element of the degree to which cultural transmission occurs and allows us to measure how important transfer of culture is in maintaining the elderly.
The consumption of middle-aged and elderly people at time $t$ are $c_{mt}$ and $c_{et}$, while current gifts to the elderly are given by $g_{mt}$. Lifetime utility of a current middle ager depends upon consumption and gifts according to the log-linear utility function:
$U_{t}=(1-\eta_{t})\ln{c}_{mt}+\eta_{t}\ln g_{t}+\beta\ln c_{et+1}$ (1)
In the lifetime utility specification (1), $\eta_{t}$ measures the relative importance of gift-giving in middle-aged utility.
Current middle-aged agents receive income $y_{mt}$ in the form of a numeraire. A middle-aged agent divides this income between consumption $c_{mt}$, savings $s_{mt}$, and gifts to the elderly $g_{mt}$. Income can also be used to teach children to appreciate culture. We model this ‘inculcative investment’ as a choice of the future value of the demand parameter $\eta_{t+1}$, that is, the value of $\eta$ that current children will have when they are middle-aged. [^16] As we noted previously, this distinguishes our model from the models of Olivera (2013) and Varvarigos (2021), which assume that inculcative investments increase the likelihood that children will have the same preferences as adults. One might refer to those models as ‘Simpsonian’ models reflecting Homer Simpson’s attitudes about parenting: “Kids are the best… You can teach them to hate the things you hate.” (The Simpsons, season 11 episode 233, 1999) The cost of setting this value in current resources is $d_{t}\eta_{t+1}$. We shall typically assume that $d_{t}=\delta y_{mt}$, where $\delta\in(0,1)$ so that inculcative costs are scaled to current income. This could be considered a simple way to capture the idea that the costs of inculcation are measured in opportunity costs, and when inculcation occurs, the value of the resources employed to inculcate is measured in units of income. [^17] Without this assumption equilibria would be non-stationary and analysis of balanced-growth paths would be impossible. Middle-aged consumption $c_{mt}$ is then income, less inculcative investment costs, savings, and gifts to the elderly, as follows.
$c_{mt}=y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}$ (2)
We now describe the components of $c_{et}$ to complete a description of lifetime utility (1). The elderly may also earn income, which we denote as $y_{et}$. $y_{et}$ might include the income earned by the elderly if they perform useful tasks or other functions, including administrative, advisory, or other types of service, as documented in Section 3. A broader interpretation of $y_{et}$ includes government policies such as social security or other government transfer payments in applications of the model to modern societies. In a hunter-gatherer society, the income $y_{et}$ received by the elderly could also derive from the productive efforts of others in the form of customary sharing of output.
Elderly consumption differs from the income of the elderly and can be increased by consumption derived from gifts received $g_{et}$. Elderly consumption may also derive from the usual source in an overlapping-generations model: returns from assets accumulated earlier in life. Elderly consumption at time $t$ then consists of returns from savings in middle age, $R_{t}s_{mt-1}$, income earned $y_{et}$, and gifts received:
$c_{et}=R_{t}s_{mt-1}+y_{et}+g_{et}.$ (3)
One point that requires clarification is how exactly this model of intergenerational gift giving relates more broadly to the transmission of culture. Shortly, we shall show that the model characterized by the utility function in (1) is observationally equivalent to a model in which cultural goods of endogenously determined value are transferred between generations.
4.1 Simple model mechanics
We begin by tackling a simplified version of the model in which there is no savings ($s_{mt}=0$ in Equation (3)). This specialization of the model serves three purposes. First, it familiarizes the reader with simple model intuition, isolating only the inculcation and gift decisions of current middle-aged agents. Second, it serves an empirical purpose in that this version of the model better characterizes premodern economies where asset formation and capital accumulation are not important, and most output depends on labor along with natural resources such as land to produce output, as is the case among hunter-gatherer societies and subsistence agricultural societies. Third, this model specialization is well suited to develop some insight into how property rights, inculcation, and elderly treatment are interrelated.
Substituting middle-aged consumption (2) and elderly consumption (3) into (1), and setting $s_{mt}=0$ gives:
$U_{mt}=(1-\eta_{t})\ln(y_{mt}-g_{mt}-d_{t})+\eta_{t}\ln g_{mt}+\beta\ln\left(y_{et+1}+g_{et+1}\right)$ (4)
To simplify things even further, we assume for the time being that $\eta_{t}\in\{0,\eta\},0<\eta<1$ so that the current middle-aged generation makes a simple binary decision about whether to inculcate children to enjoy giving gifts to the elderly at fixed intensity $\eta$ at cost $d_{t}$ or not, while at the same time choosing a gift to give to the current elderly based on the previous generation’s inculcation decision. These assumptions allow us to solve the model as one would a stationary, infinite-horizon game in which each generation sequentially decides whether or not to inculcate the youth, (correctly) anticipating that future generations will maintain the (subgame-perfect Nash equilibrium) strategy.
Middle-aged lifetime utility when inculcation has occurred in the past, is continued in the present, and is expected by current agents to continue in the future is:
$U_{mt,\eta}=(1-\eta)\ln(y_{mt}-g_{mt}-d_{t})+\eta\ln g_{mt}+\beta\ln\left(y_{et+1}+g_{et+1}\right)$ (5)
Where $g_{et+1}$ is the gift expected in the next period, given that current youth are inculcated. We can see if continuing this inculcative equilibrium is stable by contemplating what happens when a current middle-aged agent deviates and does not engage in inculcative investment. A deviation from the equilibrium path, thereby discontinuing the inculcation policy, would result in no future gifts, so the current middle-aged agent would attain lifetime utility:
$U_{mt,0}=(1-\eta)\ln(y_{mt}-g^{\prime}_{mt})+\eta\ln g^{\prime}_{mt}+\beta\ln(y_{et+1})$ (6)
In the case where gift giving and inculcation is continued, we find optimal gifts by differentiating (5) with respect to $g_{mt}$ and solving gives:
$g_{mt}^{*}=\eta(y_{mt}-d_{t})$ (7)
The current middle-aged agent therefore gives a gift proportion to disposable income, which is just income left over after inculcation. Having solved for equilibrium gifts given, we can easily solve for gifts received, $g^{*}_{et}$. As the relative size of the elderly population plays an important part in many discussions of the role of the elderly in the economy, we wish to include population changes in the model. We do this simply by assuming that the population grows (or declines) at a constant rate $n$, so $n_{t+1}=(1+n)n_{t}$. This implies that for gifts given to be equal in quantity to gifts received, so in equilibrium, $g^{*}_{et}=(1+n)g^{*}_{mt}$.
The middle-aged agent who discontinues the inculcation policy gives what amounts to a final gift of $g^{\prime}_{mt}=\eta y_{mt}$, and then receives no gift when elderly, in the following period.
4.1.1 Gifts and transfer of cultural goods
Before continuing to analyze inculcative decisions, we show that the warm glow gift-giving model is equivalent to a model in which cultural goods are transferred between generations. This allows us to relate the importance of gifting in support of the elderly to a more general notion of cultural transmission, its importance in maintaining the elderly, and its importance in society in general.
Suppose that instead of enjoying giving gifts, the current middle-aged generation places value on consumption of a cultural good, for which a price $p_{t}$ must be paid. Suppose that the demand for the cultural good is $x^{d}_{t}$ and that the good is supplied by the current elderly, each of whom has one unit of the good to supply. Lifetime utility (1) in these circumstances could be written as follows:
$U_{t}=(1-\eta_{t})\ln{c}_{mt}+\eta_{t}\ln x^{d}_{t}+\beta\ln c_{et+1}$ (8)
Setting $s_{mt}=0$ and modifying consumption of middle-aged and elderly (2) and (3) accordingly gives (8) as:
$U_{t}=(1-\eta)\ln(y_{mt}-g_{mt}-p_{t}x_{dt})+\eta\ln x_{dt}+\beta\ln\left(y_{et+1}+x_{st+1}p_{t+1}\right)$ (9)
In specification (9), $\eta$ now measures the relative importance of cultural consumption to material consumption. Current middle-aged agents spend $p_{t}x_{mt}$ on cultural goods, which gives them utility $\eta\ln x_{mt}$.
The current middle-aged inculcate an interest in cultural goods in the current young. When cultural goods are more valuable to the young, as reflected in a higher value of $\eta_{t+1}$, they pay more for them, enriching the elderly. [^18] This is similar to a model in which firms engage in advertising spending which changes the demand for a good as described in Bagwell (2007), for example. We can now establish that this cultural goods model is observationally equivalent to the warm-glow gift-giving model under our assumptions. This leads us to Proposition 1.
Proposition 1 (Equivalence of gift-giving and cultural exchange).
Given model assumptions, the models captured by utility functions (1) and (8) are observationally equivalent:
  • Gift-giving: Middle-aged agents choose an amount of gifts to give to the elderly, and receive ‘warm-glow’ utility from giving gifts according to the parameter $\eta$.
  • Cultural exchange: Middle-aged agents spend resources $x^{d}_{t}p_{t}$ on cultural goods provided by the elderly, and earn utility from these goods according to the parameter $\eta$.
Proof:.
The gift-giving model resulted in an equilibrium gift given by current middle-aged agents to the elderly that was proportional to current middle-aged disposable income, as in (7). Assuming that inculcation occurs, differentiating (9) with respect to $x^{d}_{t}$ and solving, one finds that the demand for $x^{d}_{t}$ by a current middle-aged is:
$x_{t}^{d}=\eta\frac{y_{mt}-d_{t}}{p_{t}}$ (10)
By assumption, each elderly agent has a supply of one unit of cultural goods, so demand equal to supply requires $x_{t}^{s}=(1+n)x_{t}^{d}$ or $1=(1+n)x_{t}^{d}$, as each elderly agent possesses a unit of culture. Equating demand from (10) with supply and solving gives $p_{t}=(1+n)\eta(y_{mt}-d_{t})$. From this it follows that $p_{t}x_{t}^{d}=\eta(y_{mt}-d_{t})$ is the payment middle-aged agents make in equilibrium to the elderly for cultural goods, while $p_{t+1}x^{s}_{t+1}=p_{t+1}=(1+n)\eta(y_{mt+1}-d_{t+1})$ is the amount received by the elderly, identical to the gifts implied by $g^{*}_{mt}$ in (7) and the resulting gifts received by the elderly, $g_{et}^{*}$. ∎
The equivalence result in Proposition 2 is useful in that it allows us to work with the easier-to-solve gift-giving model, but also conjecture how the importance of culture interacts with the treatment of the elderly.
4.1.2 Inculcation decisions
We can also now show that the critical thing for determining inculcation is whether or not elderly income is small relative to middle-aged income. This result and the key income threshold are described in Proposition 2.
Proposition 2 (Relative income and inculcation).
Inculcation occurs when elderly income is sufficiently small relative to middle-aged income, occurring when:
$\frac{y_{et+1}}{y_{mt+1}}\leq Y^{*}$
where
$Y^{*}=\frac{(1+n)\eta(1-\delta)}{\left((1-\delta)^{-\frac{1}{\beta}}-1\right)}$
Proof:.
Substitution of the optimal gift in (7) back into (5), and substitution of the terminal gift $g^{\prime}_{mt}=\eta y_{mt}$ into (6), coupled with the observation that $g_{et+1}=(1+n)g_{mt+1}$ allows us to write the difference in utilities along the equilibrium path with inculcation, and the utility from discontinuing inculcation, $\Delta U_{t}=U_{mt,\eta}-U_{mt,0}$ as:
$\Delta U_{t}=\ln(y_{mt}-d_{t})-\ln y_{mt}+\beta\ln(y_{et+1}+(1+n)\eta(y_{mt+1}-d_{t+1}))-\beta\ln(y_{et+1})$ (11)
$\Delta U_{t}$ in (11) simplifies to:
$\Delta U_{t}=\ln\left(1-\frac{d_{t}}{y_{mt}}\right)+\beta\ln\left(1+(1+n)\frac{\eta(y_{mt+1}-d_{t+1})}{y_{et+1}}\right)$ (12)
Using the assumption that $d_{t}$ is proportional to $y_{mt}$, so $d_{t}=\delta y_{mt}$, where $\delta\in(0,1)$ is a fixed positive constant, and substituting $d_{t}=\delta y_{mt}$ into (12) and rearranging gives:
$\Delta U_{t}\geq 0\quad\textrm{if}\quad\frac{y_{et+1}}{y_{mt+1}}\leq\frac{(1+n)\eta(1-\delta)}{\left((1-\delta)^{-\frac{1}{\beta}}-1\right)}=Y^{*}$ (13)
Proposition 2 defines a critical threshold for the relative income earned by the elderly, so that if relative elderly income falls below the threshold $Y^{*}$, inculcation occurs in equilibrium. The threshold increases (making inculcation more likely) if population growth increases ($n\uparrow$); this is because there are more middle-aged agents for each elderly agent, resulting in more gifts received. Furthermore, if the effectiveness of inculcation increases ($\eta\uparrow$), if inculcation becomes cheaper ($\delta\downarrow$), and if old-age consumption becomes more important ($\beta\uparrow$), inculcation becomes more likely. When inequality (13) does not hold, middle-agers do not anticipate the need to inculcate current youth, as the income they earn when old generates sufficient income for sustenance. At the same time, the relatively low income of middle-agers does not produce sufficiently large gifts to warrant inculcation.
We can now add to the model a measure of the status or prestige of the elderly: the relative consumption of the elderly, $\frac{c_{et}}{c_{mt}}$. Using (7) and (13), we write this consumption ratio as follows, taking into account the range of relative incomes for which inculcation occurs and substituting for equilibrium gifts when necessary:
$\frac{c_{et}}{c_{mt}}$ $=\frac{y_{et}}{y_{mt}},$ $\frac{y_{et}}{y_{mt}}>Y^{*}$ (14)
$=\frac{y_{et}}{y_{mt}(1-\eta)(1-\delta)}+\frac{(1+n)\eta}{1-\eta},$ $\frac{y_{et}}{y_{mt}}\leq Y^{*}$
Equation (14) suggests that there may not be a direct link between inculcative activity and elderly treatment. It may be the case that the elderly provide services or other productive labor of sufficient productive value so that $y_{et}$ is large relative to $y_{mt}$. The elderly will then be well treated, but this could be a result of the nature of production, not of inculcation.
As the ethnographic record emphasizes both the nature of production and different types of property rights as important determinants of the way in which the elderly interact with the economy, particularly among those societies subsisting through hunting and gathering or subsistence agriculture, we now extend the model in this direction.
4.2 Production Technology
A necessary first step in extending the analysis is the introduction of an aggregate production technology. This allows us to more fully characterize the ultimate sources of income used in expressions such as those in Proposition 2, and how they vary with the base features of the economy; for example, how labor- or land-intensive production is and even be able to distinguish between different types of labor.
Suppose that in each time period, society creates output using labor $L_{t}$ and resources that do not vary with time $T$. We will refer to $T$ as land, but it could be thought of as any other resource in addition to labor that can be owned and used to produce output. Output is produced via the constant-returns production function:
$Y_{t}=F(L_{t},T)$ (15)
Let $A_{it},i=m,e$ denote the effective labor endowments of middle-aged and elderly agents. $A_{mt}$ might be large relative to $A_{et}$ in situations where most labor is hard physical labor, as is the case among hunter-gatherers who specialize in hunting big game, or in some agricultural societies where farming is difficult and labor intensive work. On the other hand, the elderly may be able to perform some tasks that directly or indirectly provide material gains, such as providing childcare, cooking, or animal processing, which could make $A_{et}$ closer to, but still less than $A_{mt}$. We assume that all agents supply their labor inelastically, so total labor in (15) is effective labor supplies multiplied by population levels:
$L_{t}=N_{mt}A_{mt}+N_{et}A_{et},$ (16)
As the production function (15) is assumed to be constant-returns to scale, it can be decomposed using Euler’s theorem as follows:
$Y_{t}=F(L_{t},T)=F_{L}L_{t}+F_{T}T$ (17)
The decomposition in (17) gives us a natural way to conceptually attribute output to its sources. We can say $F_{L}L$ is labor’s share of output, while $F_{T}T$ is share created by fixed inputs. Similarly, we can think of the contribution of any agent’s labor to total output as $F_{L}l_{i}$, so we can think of the output created by the labor of different ages of agents as follows:
$q_{mt}^{L}=A_{mt}F_{L},\quad q_{et}^{L}=A_{et}F_{L}$ (18)
using this nomenclature, output attributable to the fixed resource is
$q_{t}^{T}=TF_{T}$ (19)
The role property rights play in the economy is to define how the output created in (18) and (19) is actually allocated to agents, as we now discuss.
4.3 Property Rights
In Section 3, we emphasized Simmons’ somewhat enigmatic statements about the importance of property rights in providing support for the elderly. Although Simmons suggests that property rights are ‘lifesavers’ for the elderly (Simmons, 1970, p. 36), he also suggests that weaker property rights in some cases are good for the elderly, for example, when food must be shared communally. Our starting point is making a distinction between property rights over inputs such as land and over output - indeed, Simmons seems to confound different conceptions of property rights in his overview of the ethnographic record. Presumably, Simmons meant to say that strong property rights over productive assets are good for the elderly, because asset ownership requires some time and resources be devoted to accumulation or establishment of ownership. But Simmons also seems to indicate that weak property rights over output (i.e., output sharing) also benefit the elderly. This distinction between property rights over output and input, we find, is addressed quite naturally in light of the Euler decomposition of the production function (17).
Let us first imagine that the amount of output created by an agent depends on how much of each productive resource is commanded by the agent. As a starting assumption, suppose that labor is inalienable and therefore any agent is able to command his or her natural share of output $q_{it}^{L}$ from (18). Also, add to this an ownership share of the fixed resource $\sigma_{it},i=m,e$ where $\sum\sigma_{it}=1$ across the population or, in our case, $\sigma_{et}N_{et}+\sigma_{mt}N_{mt}=1$. Then, a starting point to describe the share of output that an agent in some sense creates is expression (20):
$q_{it}=A_{it}F_{L}+\sigma_{it}F_{T}T,\quad i=m,e;\quad\sum_{N}\sigma_{it}=1$ (20)
With (20) as our starting point, we can describe property rights over productive resources using different configurations of $\sigma$, and we can describe property rights over output by connecting $q$, output created, to realized incomes $y_{e}$ and $y_{m}$.
4.3.1 Property rights over land
The specification in (20) naturally frames our question of the degree to which ownership rights are developed over land (or other fixed resources) in terms of $\sigma$. What values of $\sigma_{it}$ capture the idea that property rights over land are well defined and what values of $\sigma_{it}$ indicate that property rights are not well defined? Consider a situation in which ownership of land is not defined at all, that is, land is a free-access resource. $\sigma_{it}$ should then be set so as to allow our model to mimic a typical common production model, in which agent $i$’s output depends upon $i$’s relative share of labor, so $q_{it}^{L}=\frac{l_{i}}{L}F(L,T)$. [^19] This is a slight generalization of the idea that open access resources generate average product as returns for users as in, for example, Gordon (1954). See Baker and Conning (2024) for further discussion.
Therefore, to reflect the idea that land is completely open access, we use a value of $\sigma_{it}$, $\sigma_{it}^{0}$, as follows:
$\sigma^{0}_{it}=\frac{A_{it}}{L_{t}}=\frac{A_{it}}{N_{mt}A_{mt}+N_{et}A_{et}},\quad i=m,e$ (21)
To confirm that this is in accordance with a typical common property, open access model, note that we can combine agents’ output (20) with $\sigma_{it}^{0}$ in (21) to get:
$q_{it}=A_{it}F_{L}+\frac{A_{it}}{L_{t}}F_{t}T=\frac{A_{it}}{L_{t}}\left(F_{L}L_{t}+F_{t}T\right)=\frac{A_{it}}{L_{t}}F(L_{t},T)$ (22)
Where the last part of (22) follows from the assumption that $F(L_{t},T)$ is homogeneous of degree one. This pins one side of a spectrum of possible values for ownership rights over land, with $\sigma_{it}^{0}$ in (21) corresponding to a situation in which there are no land property rights.
[…]
$\sigma^{1}_{et}=\frac{1}{N_{et}},\quad\sigma_{mt}^{1}=0$ (23)
In (21) and (23), we have described free access and fully defined rights over land using values for $\sigma_{it}$ as $\sigma_{it}^{0}$ and $\sigma_{it}^{1}$. We can imagine a spectrum of intermediate possibilities between these two polar extremes and therefore introduce a parameter $\phi_{T}\in[0,1]$ to capture the degree of definition of property rights over land (or other fixed assets). This results in values the following values for $\sigma$:
$\sigma_{et}=\phi_{T}\frac{1}{N_{et}}+(1-\phi_{T})\frac{A_{et}}{L_{t}}$ (24)
$\sigma_{mt}=(1-\phi_{T})\frac{A_{mt}}{L_{t}}$ (25)
We can now say when $\phi_{T}=0$ in (24) and (25) there are no ownership rights in land, while when $\phi_{T}=1$ property rights are completely defined, with intermediate values of $\phi_{T}$ corresponding to varying degrees of ownership over land.
4.3.2 Property rights over output
We have been careful to distinguish the output $q$ actually created by an agent and the income $y$ actually received. We model imperfect rights over output as the loss of some share of $q$ to a general output fund, which is then shared equally among all agents, and do so in such a way so that we can develop expressions for $y$ for each type of agent. Let $\phi$ denote the share of own output retained by the agent so that $1-\phi$ of each agent’s output is contributed to the output pool. Then we can write the total income received by an agent as a function of the degree of definition of property rights as captured by different values of $\phi$ and $\phi_{T}$, assuming that each agent gets an equal share of the total output shared with others:
$y_{it}=\phi q_{it}+(1-\phi)\frac{\sum_{j=1}^{N_{t}}q_{jt}}{N_{t}}\quad i=m,e$ (26)
Substituting into (26) our expressions for incomes in (20) gives:
$y_{it}=\phi\left(A_{it}F_{L}+\sigma_{it}F_{T}T\right)+(1-\phi)\frac{F(L_{t},T)}{N_{t}}\quad i=m,e$ (27)
In (27) we see that imperfect ownership of output removes a bit of output income from the agent but entitles an agent to a share of average output.
4.3.3 Property rights in general
A final step allows us to discuss the development of property rights in a unified way, rather than discussing property rights in output or inputs. [^21] Developing a theory of the development of property rights is beyond the scope of this paper, but there is no shortage of papers discussing this. For a recent review and one modeling approach, see Baker and Conning (2024). We are supposing that the development of property rights is exogenous in this paper, but typically it is the case that more developed economies have more private property rights. See, e.g., Baker (2008) for a model in which rights over assets emerge endogenously, and Baker and Swope (2021) for a model in which ownership of resources and output-sharing are interrelated. A stylized fact is that property rights tend to develop in tandem; ownership of resources and the output created by those resources seem to evolve in the same direction. We model this fact by assuming that $\phi_{T}$ and $\phi$ are linked. To allow this linkage to be flexible and for the degree of different kinds of property rights to develop at different rates, we introduce $\rho$ as a parameter characterizing the relative speed at which ownership of output and inputs co-evolve and then specify that $\phi_{T}=\phi^{\rho}$.
Since $\phi\in[0,1]$ it follows that when $\rho>1$, $\phi_{T}=\phi^{\rho}<\phi$, and one might say that rights over output are more developed than rights over assets for any interior value of $\phi$. Alternatively, if $\rho<1$, the opposite is true: rights over assets are more developed than rights over output, as $\phi_{T}>\phi$. This approach allows us to acknowledge that different kinds of property rights develop together, while also allowing us to remain agnostic as to which kinds of rights develop faster.
4.3.4 Back to relative treatment of the elderly
We now can relate the development of property rights to inculcation, treatment of the elderly, and cultural transmission, which means describing how property rights relate to the results in Propositions 1 and 2. Substituting the definitions (24) and (25) into (27), and forming the ratio $\frac{y_{et}}{y_{mt}}$ gives:
$\frac{y_{et}}{y_{mt}}=\frac{\phi\left(A_{et}F_{L}+\left[\frac{A_{et}(1-\phi^{\rho})}{L_{t}}+\frac{\phi^{\rho}}{N_{et}}\right]F_{T}T\right)+(1-\phi)\frac{F(L,T)}{N_{et}+N_{mt}}}{\phi\left(A_{mt}F_{L}+\frac{A_{mt}(1-\phi^{\rho})}{L_{t}}F_{T}T\right)+(1-\phi)\frac{F(L,T)}{N_{et}+N_{mt}}}$ (28)
Assume that the aggregate production function is Cobb-Douglas: $F(L,T)=L^{\alpha}T^{1-\alpha}$. Then, expression (28) simplifies to:
$\frac{y_{et}}{y_{mt}}=\frac{\phi\left(A_{e}\alpha+\left[A_{e}(1-\phi^{\rho})+\frac{\phi^{\rho}L_{t}}{N_{et}}\right](1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}{\phi\left(A_{mt}\alpha+A_{mt}(1-\phi^{\rho})(1-\alpha)\right)+(1-\phi)\frac{L_{t}}{N_{et}+N_{mt}}}$ (29)
Expression (29) links relative elderly income, technology of production, and property rights. It provides direct insight into how well the elderly are treated given the fundamental characteristics of the economy on its own, but it also functions as the driving variable in whether or not we expect to see inculcation and a cultural factor in the treatment of the elderly as it appears in the inequality (13) in Proposition 2.
When $\phi=0$, (29) collapses to unity, as one would expect. This is because when all output is shared equally, elderly and middle-aged are the same. In Proposition 3, we collect and further define some of the key results concerning relative income and the definition of property rights.
Proposition 3 (Property rights and the relative income of the elderly).
The relative income of the elderly, $\frac{y_{et}}{y_{m}t}$ is U-shaped in the degree of definition of property rights $\phi$, provided a) production is not too labor-intensive ($\alpha$ is not too large), and provided $A_{e}<A_{m}$; the elderly have a relatively smaller endowment of labor.
Proof:.
In what follows, the total population is $N_{et}+N_{mt}=N_{t}$. Taking the derivative of relative income $\frac{y_{et}}{y_{mt}}$ in (29) and simplifying gives:
$\frac{\partial\left(\frac{y_{et}}{y_{mt}}\right)}{\partial\phi}=\frac{N_{t}L_{t}\left(A_{et}N_{et}+A_{mt}(N_{t}\phi^{\rho}(1-\alpha)(1+\rho)-N_{et})\right)}{N_{et}\left((1-\alpha)A_{mt}N_{t}\phi^{\rho+1}-(A_{mt}-A_{et})N_{et}\phi-L_{t}\right)^{2}}$ (30)
Evaluating (30) at $\phi=0$ gives:
$\frac{\partial\left(\frac{y_{et}}{y_{mt}}\right)}{\partial\phi}=(A_{et}-A_{mt})\frac{N_{t}}{L_{t}}$ (31)
Which is negative so long as $A_{et}<A_{mt}$. Therefore, the relative income ratio starts at $1$ and initially decreases in $\phi$. The sign of the derivative in (30) thereafter depends wholly on the expression in the numerator of (30):
$A_{et}N_{et}+A_{mt}(N_{t}\phi^{\rho}(1-\alpha)(1+\rho)-N_{et})$ (32)
This expression is increasing in $\phi$ and equals zero at a critical value of $\phi$, $\phi^{*}$, given by:
$\phi^{*}=\left(\frac{(A_{mt}-A_{et})N_{t}}{A_{mt}N_{t}(1-\alpha)(1+\rho)}\right)^{\frac{1}{\rho}}$ (33)
As the derivative is initially negative, equals zero at a unique point $\phi^{*}\in(0,1)$ under the maintained assumptions, and then becomes positive, it follows that relative income is U-shaped in the degree to which property rights are developed, $\phi$. ∎
The proposition starts with the assumption that $A_{et}<A_{mt}$, which means that an elderly agent provides fewer effective units of labor than a middle-aged agent. If this assumption did not hold, the relative income of the elderly would be strictly increasing in $\phi$. Moreover, if $\alpha$ were very large, $\phi^{*}$ would be outside the $(0,1)$ range, so relative elderly income only decreases in $\phi$. Note that when $\phi=1$ we have:
$\frac{y_{et}}{y_{mt}}=\frac{1}{\alpha}\frac{A_{et}}{A_{mt}}+\frac{1-\alpha}{\alpha}\frac{N_{mt}}{N_{et}}$ (34)
The limit ratio in (34) can be greater than one or less than one, depending on the relative elderly population and their relative productivity. For small enough $\alpha$ (that is, production is resource intensive), the ratio in (34) can be greater than one even if $A_{et}=0$ (when the elderly do not supply any material labor to the economy). In these circumstances, the elderly can achieve a high level of prestige measured by relative income even in the absence of inculcation. This is because the elderly have accumulated crucial resources while in middle age.
Figure 1: Relative consumption of the elderly ($\frac{c_{et}}{c_{mt}}$), as functions of the development of property rights in output and productive resources from nonexistent ($\phi=0$) to complete ($\phi=1$). Baseline parameter values of $A_{m}=1$, $A_{e}=0.025$, and $\alpha=0.5$.
Figure 2: Relative consumption of the elderly ($\frac{c_{et}}{c_{mt}}$), as a function of the degree to which property rights are developed, $\phi$, with endogenous inculcation. Baseline parameter values of $A_{m}=1$, $A_{e}=0.025$, and $\alpha=0.5$.
In Figure 1, the relative income ratio (29) is plotted as a function of the degree of development of property rights, $\phi$, for different assumptions about the underlying nature of production. Although changes in model parameters influence its exact position, it retains the same overall shape. The figure is suggestive of the empirical evidence discussed in the previous section. For one, it shows how, possibly, in a relatively egalitarian society, there is no need for cultural protection of the elderly, but as property rights develop, at least initially, a demand for some other means of ensuring the relative livelihood of the elderly arises as relative consumption of the elderly experiences downward pressure, and developing culture is a way to accomplish this. The provision of cultural goods and things like food taboos is a way of transferring current income for future consumption via inculcation. However, as the possibilities for the ownership of resources become more expansive, this helps the elderly as they are able to benefit from the income that owned assets provide. This may dull the need for direct institutions, such as food taboos, while at the same time the prestige of the elderly may continue to increase.
In Figure 2, we illustrate some possibilities graphically. The figure essentially combines the results of Propositions 2 and 3. As property rights develop ($\phi$ grows larger) and the economy moves to the right along the $x$ axis, $\frac{c_{et}}{c_{mt}}$, which is equivalent to $\frac{y_{et}}{y_{mt}}$ in this range, initially falls and passes the $Y^{*}$ threshold in Proposition (2) at the point marked by the vertical dashed line, so inculcation of young people to give gifts to the elderly begins. This mitigates the decline in the relative treatment of the elderly in the indicated range. However, inculcation ceases once property rights become sufficiently developed, as indicated in the figure.
This analysis suggests that it might be difficult to forge a direct link between culture, inculcation, and the prestige of the elderly, and that any relationship between culture, the means of production, and elderly treatment may be nonlinear.
4.4 Savings, Capital Accumulation, and Inculcation
We now analyze a version of the model in which assets are accumulated via conscious savings decisions, and investigate how accumulation interacts with inculcation and gifts to the elderly. This allows us to study how inculcative investment interacts with savings and also allows us to develop the exposition in a way that is relatable to classical growth models with technological progress.
The optimization problem of a current middle-aged agent can be written as the problem of maximizing $U_{t}$ in (1) subject to constraints (2) and (3), as follows:
$\max_{g_{t},s_{mt},\eta_{t+1}}$ $(1-\eta_{t})\ln{c}_{mt}+\eta_{t}\ln g_{mt}+\beta\ln c_{et+1}$ (35)
s.t. $\quad c_{mt}=y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt},$
$\quad c_{et+1}=s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}$
The solution of (35) presents two challenges. First, a solution must take into account that current middle-aged agents’ future gifts depend on current inculcative investment through the parameter $\eta_{t+1}$, and in an equilibrium solution to the model, beliefs about this dependence must be correct and consistent with current actions. Second, a solution must also admit the possibility of corner solutions in which $\eta^{*}_{t}=0$. The subgame-perfect Nash equilibrium solution to the problem (35) has a structure in which agents give gifts in proportion to the present discounted value of lifetime income, with this proportion depending on $\eta_{t}$. For the time being, we have bypassed the issue of the determination of $\eta_{t}^{*}$ itself; as an equilibrium object, its determination can only be addressed in a fully specified general equilibrium setting. The exact form of the equilibrium gift-giving function is given in Proposition 4:
Proposition 4 (Gifts with accumulation).
If $\eta_{t}>0$ in equilibrium, the $g^{*}_{mt}$ is given by:
$g^{*}_{mt}=\frac{\eta_{t}}{1+\beta}\left(y_{mt}+\frac{y_{et+1}}{R_{t+1}}\right)$ (36)
Proof:.
See Appendix A. ∎
Proposition 4 shows how equilibrium gifts and therefore, by analogy to Proposition 1, cultural transmission, are proportional to lifetime income. Using Proposition 4, we can then begin to solve the maximization problem in (35). Two aspects of this solution are highlighted in Proposition 5.
Proposition 5 (Total savings and gross returns on savings).
If $\eta_{t}>0$ in equilibrium, middle-aged agents’ aggregate savings - the sum of inculcative investments and conventional savings - is described by:
$s_{mt}+\eta_{t+1}d_{t}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (37)
In equilibrium, the gross rate of return on conventional savings is given by:
$R_{t}=\frac{1+n}{1+\beta}\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{d_{t-1}}$ (38)
Proof:.
See Appendix B. ∎
Equation (37) in Proposition 5 describes the total amount of current income middle-aged agents wish to allocate to future consumption. This condition shows that expenditures on savings and inculcation add up to the fraction of income a middle-aged agent wishes to save, which depends upon current income less discounted future income. Equation (38) in Proposition 5 requires that the rates of return on the different means of transferring current wealth to the future in (37) be equilibrated. That is, the rate of return on inculcative investments and what can be called conventional savings must be equal in equilibrium; otherwise, middle-aged agents would not employ both savings instruments. As we maintain the assumption that $d_{t}$ is proportional to income ($d_{t}=\delta y_{mt}$), we can interpret (38) to mean that returns to accumulated assets such as capital must be equal in equilibrium to the unit return on inculcative investment, which weighs future gifts against current inculcative costs.
The question remains whether an equilibrium with gifts and inculcation is sustainable, as both Propositions 4 and 5 do not allow much further analysis of the nature and magnitude of inculcative investments, and also proceed following the assumption that $\eta_{t}>0$. To go further, we need to round out the model to see how $\eta_{t}$, $s_{t}$, and $R_{t}$ depend on the underlying nature of the economy, as they are quantities determined in equilibrium and require specification of the production side of the economy.
4.4.1 Production and Economic Growth
To fully study the model, we need to specify the underlying nature of production in the economy. This will allow us to turn the two equations in Proposition 5 into dynamic laws describing how the effective capital stock progresses and how inculcative investments proceed. We again use a simple Cobb-Douglas production technology in which production occurs according to the aggregate production function $Y_{t}=L_{t}^{\alpha}K_{t}^{1-\alpha}$, where $K_{t}$ now denotes the total capital stock and as before $L_{t}$ denotes total effective labor supply:
$L_{t}=A_{mt}N_{mt}+A_{et}N_{et}$ (39)
We assume that $A_{it},i=e,m$ grows at a constant exogenous rate $a$, while the population grows at a rate $n$. As in previous sections, constant population growth means that $N_{mt}=(1+n)N_{et}$, and that $\tfrac{1}{1+n}$ is a measure of the relative size of the elderly population.
We use (39) and the production technology to obtain the following expressions for middle-aged labor income, elderly labor income, and the gross rate of return on capital, under the maintained assumption that all resources are paid their marginal products, where $k_{t}=\frac{K_{t}}{L_{t}}$ is capital per effective unit of labor:
$y_{mt}$ $=$ $\alpha A_{mt}\left(\frac{K_{t}}{L_{t}}\right)^{1-\alpha}$ $=$ $\alpha A_{mt}k_{t}^{1-\alpha}$ (40)
$y_{et}$ $=$ $\alpha A_{et}\left(\frac{K_{t}}{L_{t}}\right)^{1-\alpha}$ $=$ $\alpha A_{et}k_{t}^{1-\alpha}$
$R_{t}$ $=$ $(1-\alpha)\left(\frac{K_{t}}{L_{t}}\right)^{-\alpha}$ $=$ $(1-\alpha)k_{t}^{-\alpha}$
Applying these equations to the equilibrium relationships (37) and (38) in Proposition 5 admits the description of the equilibrium paths of $\eta$ and $k$. Pursuant to analyzing (40), we also define some new terms:
[…]
Given our assumptions of a constant rate of technological progress, we can omit time subscripts for the terms introduced in (41) and write $\tau_{e}$ and $\gamma_{e}$. One can also show that $\tau_{e}$ and $\gamma_{e}$ are related via $\gamma_{e}=\frac{\tau_{e}}{\tau_{e}+(1+a)}$.
Application of factor returns in (40) to (37) and (38) yields Proposition 6, which translates Proposition 5 into a pair of dynamic equations that describe the evolution of the capital stock and inculcation:
Proposition 6 (Dynamics of inculcation and savings).
In equilibrium, Proposition 5 implies that the effective capital stock evolves over time according to:
$(1-\alpha)k_{t}^{-\alpha}=\frac{(1+n)(1+a)}{(1+\beta)\delta}\left(\left(\frac{k_{t}}{k_{t-1}}\right)^{1-\alpha}+\frac{\tau_{e}(1+a)}{1-\alpha}\frac{k_{t+1}}{k_{t-1}^{1-\alpha}}\right)$ (42)
while the inculcative parameter $\eta_{t}$ evolves over time according to:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{k_{t+1}}{k_{t}^{1-\alpha}}\left(\frac{\tau_{e}(1+a)}{(1-\alpha)(1+\beta)\delta}+\frac{(1+n)(1+a)}{\alpha\delta(1-\gamma_{e})}\right)$ (43)
Proof: .
See Appendix C. ∎
Equation (43) describes the dynamics of inculcation but can also be viewed as a condition that describes whether or not there will be any inculcative investment at all, and therefore allows us to delineate when inculcative equilibria will be sustainable. When the right-hand side of (43) is nonpositive, $\eta_{t+1}=0$ in equilibrium. One can see that this possibility depends on the rates of population growth and technological progress; in particular, a higher rate of technological progress makes $\eta=0$ more likely. This result occurs because higher growth rates increase the demand and hence returns on capital, making inculcation less desirable. Also making $\eta=0$ more likely in equilibrium are higher inculcative costs, as measured by $\delta$, and greater direct participation in the economy by the elderly (larger values for both $\tau_{e}$ and $\gamma_{e}$).
A value of the relative consumption ratio of the elderly $\frac{c_{et}}{c_{mt}}$ can be found using the equilibrium values of savings, inculcation, and gifts. In Appendix D, we describe the calculations that lead to Proposition 7:
Proposition 7 (Relative consumption, inculcation, and growth).
In equilibrium, when $\eta_{t}^{*}>0$, the relative ratio of elderly to middle-age consumption is:
$\frac{c_{et}}{c_{mt}}=\frac{\beta R_{t}}{1-\eta_{t}}\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}$ (44)
Otherwise, relative consumption is simply determined by relative incomes; that is
$\frac{c_{et}}{c_{mt}}=\frac{y_{et}+R_{t}s_{t-1}}{y_{mt}}$ (45)
Proof: .
See Appendix D. ∎
Equation (44) in Proposition 7 indicates that the relative consumption ratio depends on relative lifetime incomes, and therefore the overall growth rate of the economy. The expression also suggests that many of the main general equilibrium implications of the model will flow through equilibrium asset returns, $R$, and the magnitude of $\eta$. This is particularly true along a balanced growth path, as in a balanced growth equilibrium, incomes grow at a rate $a$, and asset returns are constant. Along a balanced growth path, (44) reduces to:
$\frac{c_{e}}{c_{m}}=\frac{\beta R}{(1-\eta)(1+\alpha)}$ (46)
Equation (46) therefore indicates that parameters such as $\alpha$, the labor intensity of production, $n$, the growth rate of population, and $\delta$, the per-unit-income costs of inculcation will influence the relative consumption of the elderly by exerting an impact on the level of inculcation $\eta$ and equilibrium asset returns $R$. Of course, $\eta$ and $R$ are somewhat complicated functions of model parameters. We will postpone discussion of the complete balanced-growth solution of the model until Section 4.5.2.
4.5 Analysis with accumulation
The rest of the modeling section of the paper consists in delving into the implications of Propositions 5, 6, and 7. We first solve a simplified version of the model in which all elderly income derives from asset accumulation, and $y_{e}=0$ at every $t$, which is equivalent to assuming that $\tau_{e}=\gamma_{e}=0$, so the elderly do not contribute any productive labor and only provide previously accumulated capital to the economy.
4.5.1 Asset income as the sole source of elderly income
When $y_{et}=0$, the equilibrium gift function (36) reduces to:
$g_{mt}^{*}=\frac{\eta_{t}}{1+\beta}y_{mt}$ (47)
And the rate equilibration condition (38) is:
$R_{t}=\frac{1+n}{1+\beta}\frac{y_{mt}}{\delta y_{mt-1}}$ (48)
The sum of cultural investment and asset savings (37) becomes:
$s_{mt}+\eta_{t+1}d_{t}=\frac{\beta}{1+\beta}y_{mt}$ (49)
Equations (47, 48 and 49) reflect that when there is no elderly labor income, the decisions of current middle-ageds reduce to allocating shares of income between overall savings, gift-giving and inculcation, in such a way that the returns between inculcative investment and asset accumulation are equilibrated.
The general equilibrium aspects of the model can be deduced by substituting $\tau_{e}=0$ into (42) in Proposition 6, which gives a simple recursion for the stock of capital per effective unit labor:
$k_{t+1}=\frac{(1-\alpha)(1+\beta)\delta}{(1+n)(1+a)}k_{t}^{1-\alpha}$ (50)
Using $\gamma_{e}=0$ and $\tau_{e}=0$ in (43) in Proposition 6 and rearranging results in another recursion for the effective capital stock:
$k_{t+1}=\frac{\alpha}{(1+a)(1+n)}\left(\frac{\beta}{1+\beta}-\delta\eta_{t+1}\right)k_{t}^{1-\alpha}$ (51)
Equation (51) shows how the possibility of inculcation and gifts impacts the canonical overlapping-generations model with population and technological progress, as developed in, for example, Acemoglu (2009). But for the $\delta\eta_{t+1}$ term in (51), the accumulation condition is a standard capital accumulation condition in growth models, so the possibility of inculcative investment reduces the amount of capital per effective unit of labor, which will increase equilibrium asset return rates. Dividing (50) by (51) and solving the result for $\eta_{t+1}$ gives a time-independent solution for the inculcative parameter $\eta$:
$\eta_{t+1}^{*}=\eta^{*}=\frac{\beta}{(1+\beta)\delta}-\frac{1-\alpha}{\alpha}(1+\beta)$ (52)
Equation (52) indicates that a larger $\delta$, not surprisingly, reduces the equilibrium size of $\eta^{*}$. An increase in the relative intensity of capital in production, measured by $\frac{1-\alpha}{\alpha}$, also negatively impacts $\eta$. Societies that employ more capital-intensive means of production, other things equal, will have lower degrees of inculcation and perhaps transmission of culture.
If parameters are configured so that $\eta^{*}\leq 0$, in equilibrium $\eta=0$, and there is no inculcative investment. The elderly ensure their livelihood as they would in the typical overlapping generations model: through accumulation of productive assets during middle age. Setting (52) to zero and solving for $\alpha$ gives an interpretation of (52) as a condition for inculcation to occur in terms of the relative capital intensity of production:
$\frac{1-\alpha}{\alpha}\leq\frac{\beta}{\delta(1+\beta)^{2}}$ (53)
This critical threshold depends positively on the costs of inculcation, and on $\beta$, which influences the relative weight placed on consumption in old age in utility. If (53) is not satisfied and no inculcation occurs, we recover the law of motion for the capital stock by using $\eta=0$ in (51). In forming the relative ratio of consumption, we find it most straightforward to solve the model in terms of $R$, equilibrium asset returns, rather than in terms of $k$, the equilibrium per-unit-labor capital stock. We find a steady-state value of $k$, $\overline{k}$, and then use this to compute the balanced-growth rate of asset returns $\overline{R}=(1-\alpha)\overline{k}^{-\alpha}$.
When no inculcation occurs (that is, (53) is not satisfied), equilibrium returns can be found by using $\eta=0$ in (51). Writing the resultant expression in terms of $\overline{R}$ gives:
$\overline{R}=(1-\alpha)\frac{(1+a)(1+n)(1+\beta)}{\beta\alpha}$ (54)
When (53) is satisfied, we find $\eta^{*}$ as given in (52), and equilibrium asset returns follow from (51). These are:
$\overline{R}=\frac{(1+a)(1+n)}{(1+\beta)\delta}$ (55)
Collecting all of these results gives us a relative consumption ratio that depends on whether or not inculcation and gift giving will occur in equilibrium, which can be obtained by using equilibrium values of $\overline{R}$ and $\eta$ in Equation (46) in Proposition 7:
$\frac{c_{e}}{c_{m}}=$ $\frac{1-\alpha}{\alpha}(1+\beta)(1+n),$ $\eta^{*}=0$ (56)
$\frac{\beta(1+n)}{\delta(1+\beta)(1+\frac{1-\alpha}{\alpha}(1+\beta))-\beta},$ $\eta^{*}>0$
Increases in the rate of population growth increase the relative well-being of the elderly, whether or not there are gifts involved. In the high-capital intensity world, in which there is no gift-giving and inculcation, it is because population growth increases the relative scarcity of capital, which increases rates of return on accumulated assets, benefiting the elderly. This is also true in the second case, but the effect is augmented by another aspect of how gifts interact with population growth in the model - when there are more young people relative to elderly, there are more gifts to go around.
The relative intensity of capital in production, $\frac{1-\alpha}{\alpha}$, plays a key role in influencing the relative treatment of the elderly. The underlying reasons are similar to those described in the context of the model specification developed in Section 4.2. When production is more labor-intensive, the relative treatment of the elderly tends to decline, but at some point, inculcation and acculturation of elderly treatment come into play, mitigating this trend. Put another way, when income depends mainly on labor and not accumulated assets, there is an incentive to develop cultural practices facilitating transmission of income to the elderly, as there are no other ways to accumulate. The flip side is that when production is more asset-intensive, the relative treatment of the elderly may improve while simultaneously acculturation to treat the elderly well wanes. The full ramifications of the elderly treatment model are shown in Figure 3, which shows the relative consumption of the elderly $\frac{c_{e}}{c_{m}}$ in (56) as a function of the capital intensity of production. One can see from the figure the same sort of emergent nonlinear, U-shaped arrangement between relative elderly treatment as occurred before in the property-rights-based model of Section 4.2.
Figure 3: Relative consumption of the elderly ($\frac{c_{et}}{c_{mt}}$), as a function of the capital intensity of production, $\frac{1-\alpha}{\alpha}$.
4.5.2 Full model solution
The results of the previous Section applied when the sole source of income for the elderly is previously accumulated assets. When the elderly can save, but also have other sources of income deriving from their labor available, the resulting model is more complex because not only does elderly labor income provide sustenance when elderly, but it also raises lifetime income, which influences incentives both to give gifts and to preserve wealth for future consumption. These effects can make it difficult to accurately state how changes in model parameters influence inculcation, accumulation, and relative treatment of the elderly. In spite of this, one can still solve the model assuming that all model variables are along a balanced growth path. With $\tau_{e},\gamma_{e}>0$, the steady-state relationships corresponding to (42) and (43) in Proposition 6 can be written as:
$R=\frac{(1+n)(1+a)}{(1+\beta)\delta}\left(1+\frac{(1+a)\tau_{e}}{R}\right)$ (57)
with
$\eta=\frac{\beta}{\delta(1+\beta)}-\frac{(1+a)}{\delta R}\left(\frac{\tau_{e}}{1+\beta}+\frac{(1+n)}{(1-\gamma_{e})}\frac{1-\alpha}{\alpha}\right)$ (58)
Solving (57) for the equilibrium asset return rate gives:
$R=\frac{(1+n)(1+a)}{2\delta(1+\beta)}\left(1+\sqrt{1+\frac{4(1+\beta)\delta}{(1+n)}r_{e}}\right)$ (59)
From (57), we see that as $\tau_{e}$ increases, meaning that the elderly have a relatively larger source of income, $R$ increases in an equilibrium with a positive amount of inculcation ($\eta>0$). This is because when a future source of income is available, the incentives to save are deadened. For the same reason, increases in $\tau_{e}$ lessen inculcative investments, as is evident from (58).
When there is no inculcation, we can solve (43) with $\eta=0$ and rearrange terms to get the following expression for asset returns:
$R=\frac{1+a}{\beta}\left(\frac{(1+n)(1+\beta)}{1-\gamma_{e}}\frac{1-\alpha}{\alpha}+\tau_{e}\right)$ (60)
Using (60) and $\eta=0$ in (46), we find that relative consumption becomes.
$\frac{c_{e}}{c_{m}}=\frac{(1+n)(1+\beta)}{1-\gamma_{e}}\frac{1-\alpha}{\alpha}+\tau_{e}$ (61)
We can rearrange expression (46) to give a condition for inculcation to occur once again. This condition again reflects the key role played by the capital intensity of production. No inculcation and gift-giving occur in equilibrium so long as:
$\frac{1-\alpha}{\alpha}\geq\left(\frac{R\beta}{(1+a)}-\tau_{e}\right)\frac{1-\gamma_{e}}{(1+\beta)(1+n)}$ (62)
Figure 4: Relative consumption of the elderly ($\frac{c_{et}}{c_{mt}}$), as a function of relative capital intensity with $\tau_{e}>0$. The solid lines show how the rate of return on capital is constant while $\eta$ is positive. When capital intensity reaches a certain threshold, inculcation ends and the rate of returns starts rising. The dashed lines show how these relationships change when the elderly provide productive labor $\tau_{e}>0$.
It is easiest to see how equilibrium is influenced when $\tau_{e}>0$ compared to the case in which $\tau_{e}=0$. as depicted on Figure 3. In Figure 4, we superimpose the critical line depicting the relative capital intensity at which $\eta$ goes to zero, and the resulting consumption ratios when $\tau_{e}>0$ on those depicted in 3.
Although Figure 4 suggests greater elderly labor income reduces the incidence of inculcative, gift-giving equilibria, it also indicates that the impact of labor income on relative elderly consumption are not obvious. On the left-hand side of the figure we see that the relative consumption ratio of the elderly goes down as a result of the increased income shift. This is because total labor supply increases, which depresses labor incomes in general.
This last set of results indicates how the advent or expansion of societal welfare targeted towards the elderly may impact the transmission and maintenance of culture, some of the particulars of which we described in Section 3. Such changes act to increase $\tau_{e}$ in the model, which may have a deleterious effect on inculcation, and therefore cultural transmission.
5 Discussion and Conclusions
The model presented in this paper serves to reinforce previous work on the relationship between intergenerational transfers and their role in delivering support to the elderly. But our model also allows us to discuss more thoroughly the features of the extensive cross-cultural research that has been done on the treatment of the elderly. Our analysis spans not only relatively recent phenomena such as the demographic transition, but also helps explain earlier changes in the treatment of the elderly across the entire spectrum of levels of development. Our model extends to matters such as the development of property rights and how property rights interact with culture and elderly treatment. Our model focuses attention squarely on elderly treatment, not on intermediate features of the economy such as the occurrence and direction of transfers.
Our model demonstrates that the relationship between things such as ownership, accumulation, and production and the treatment of the elderly is intrinsically complex, even as we are able to lend theoretical support for features of the ethnographic record such as the ‘curvilinear hypothesis.’ Moreover, with our model we are also able to discuss how the importance of accumulation in production impacts the elderly. Our findings here are also subtle; the possibility of accumulation may initially have an adverse impact on the relative status of the elderly. However, increasing the importance of accumulation may eventually function to the relative advantage of the elderly. Indeed, our results lend further support to the hypothesis that there is a nonlinear relationship between the complexity of familial structure and development (Nimkoff and Middleton, 1960; Blumberg and Winch, 1972). This observation in fact echoes the general trend observed by (Murdock and Provost, 1973) - that the most culturally sophisticated societies tend to lie at the middle spectrum of development, who go so far to say that “This essentially bimodal or curvilinear distribution is inconsistent with any unilinear interpretation of social development,” (Murdock and Provost, 1973, p. 392) our model suggests why such a pattern might exist, at least in regards elderly treatment and coincident transmission of culture. [^22] This tendency has been described in the context of land inheritance rules by Baker and Miceli (2005) and in marriage institutions (Baker and Jacobsen, 2007a, b).
There are several possible avenues for further research. Some are suggested by our theory, while others derive from aspects of the ethnographic record that we did not treat in detail for reasons of brevity. One such topic is the relationship between family structure, technical change, and the treatment of the elderly. To the extent that gift giving can imply a closer family, our model informs some of these issues, but they remain an interesting avenue for future research, along with the relationship between elderly treatment, marriage, and fertility. [^23] These issues are partially explored in Cigno et al. (2020) and in Bisin and Verdier (2000). We also have not addressed gender differences in elderly treatment, and to be sure, documenting such differences is a point of emphasis of Simmons (1970). There are also potential complexities of the problem of inculcation and cultural transmission more generally that could be further explored. To the extent that inculcation and elderly treatment are behavioral norms, there may be important dynamic and public-goods aspects to decisions that govern these phenomena, some of the complexities of which figure prominently in Bowles (1998). Our model has also focused only on inter-societal relative well-being of the elderly, but a remaining issue to be explored is inequality among the elderly, where poverty differences and inequality among the elderly may be related to cultural transmission and gift giving as well. Our hope is that our general approach will be fruitful when applied to these additional topics.
Appendix A Proof of Proposition 4
We describe the equilibrium solution to the maximization problem (35). Our general plan of attack is to proceed as if the solution to the problem is interior and then to find this solution. The solution then allows us to check when it is in fact interior. We start by assuming that agents maintain the belief that in equilibrium gifts are some multiplicative function of the value of $\eta_{t}$ and model parameters as follows:
$g_{mt+1}=\eta_{t+1}\Psi_{t+1}$ (A.1)
where $\Psi_{t+1}$ is to be determined. Substituting for $g_{mt+1}$ using (A.1) into the middle-aged agent’s optimization problem (35), we find that the first-order conditions associated with an interior solution are:
$-\frac{1-\eta_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}}+\frac{\eta_{t}}{g_{mt}}$ $=0$ (A.2)
$-\frac{1-\eta_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}}+\beta\frac{R_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$
$-\frac{(1-\eta_{t})d_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}-g_{mt}}+\beta\frac{(1+n)\Psi_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$
The first expression in (A.2) implies that gifts are proportional to income, after inculcative expenditures and savings are subtracted out:
$g_{mt}=\eta_{t}(y_{mt}-d_{t}\eta_{t+1}-s_{mt})$ (A.3)
Substituting $g_{mt}$ from (A.3) back into the second and third equations in (A.2) gives two conditions describing savings and inculcation behavior:
$-\frac{1}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}}+\beta\frac{R_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$ (A.4)
$-\frac{d_{t}}{y_{mt}-d_{t}\eta_{t+1}-s_{mt}}+\beta\frac{(1+n)\Psi_{t+1}}{s_{mt}R_{t+1}+(1+n)g_{et+1}+y_{et+1}}$ $=0$
The two first-order conditions in (A.4) imply that for both savings and inculcation to occur, it must be the case that:
$\Psi_{t+1}=\frac{d_{t}R_{t+1}}{1+n}$ (A.5)
Multiplying both sides of (A.5) by $\eta_{t+1}$ then indicates that $g_{mt+1}$ in A.1 can be written as:
$g_{mt+1}=\eta_{t+1}\frac{d_{t}R_{t+1}}{(1+n)}$ (A.6)
Since $g_{et+1}=(1+n)g_{mt+1}$, we find that $g_{et+1}=\eta_{t+1}d_{t}R_{t+1}$. Using this in either part of (A.4) gives an expression describing how total savings, in effect, are distributed between asset accumulation $s_{mt}$ and inculcation $\eta_{t+1}$, which is reported in the text as equation (37):
$s_{mt}+\eta_{t+1}d_{t}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (A.7)
Expression (A.7) says that the typical middle-aged agent desires to save a fraction of current income for future savings. This savings is divided between actual savings, $s_{mt}$, and inculcative cultural investments, $\eta_{t+1}d_{t}$. Savings is also reduced according to the current discounted value of future income earned when elderly. The equilibrium condition (A.5) implies that if an agent is to engage in both asset accumulation and inculcation, the returns to each activity must be equal. Inserting the right-hand side of (A.7) into the right-hand side of (A.3) gives the equilibrium gift function in the text, equation (36):
$g^{*}_{mt}=\frac{\eta_{t}}{1+\beta}\left(y_{mt}+\frac{y_{et+1}}{R_{t+1}}\right)$ (A.8)
Comparing (A.1) and (A.8), we see that
$\Psi_{t}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (A.9)
Appendix B Proof of Proposition 5
The first part of the proposition concerning aggregate savings follows directly from Appendix A, and, in particular, Equation (A.7). To get the second part of the Proposition concerning returns on savings, iterate (A.5) back one period and substitute in the value of $\Psi_{t}$ so obtained into (A.9) to get the rate of return equalisation condition:
$R_{t}=\left(\frac{1+n}{1+\beta}\right)\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{d_{t-1}}$ (B.10)
Which (B.10) appears in the text as (38).
Appendix C Proof of Proposition 6
We begin by rewriting (37) as follows, using the assumption that $d_{t}=\delta y_{mt}$:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{y_{et+1}}{\delta y_{mt}}\frac{1}{R_{t+1}(1+\beta)}-\frac{s_{mt}}{\delta y_{mt}}$ (C.1)
Substitution for $R_{t+1}$, $y_{mt}$ and $y_{et+1}$ using (40) in (C.1), and also noting that the future capital stock depends on savings by middle-aged agents in the present, $K_{t+1}=s_{mt}N_{mt}$, gives:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et+1}k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{N_{mt}\delta\alpha A_{mt}k_{t}^{1-\alpha}}$ (C.2)
Multiplying the last term on the right by $L_{t+1}/L_{t+1}$, and substituting $N_{mt+1}=(1+n)N_{mt}$, $A_{mt+1}=(1+a)A_{mt}$, and $A_{et+1}=(1+a)A_{et}$ into (C.2), we have:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{A_{et}(1+a)k_{t+1}}{(1-\alpha)(1+\beta)\delta A_{mt}k_{t}^{1-\alpha}}-\frac{K_{t+1}}{L_{t+1}}\frac{(1+a)(1+n)}{\frac{N_{mt+1}A_{mt+1}}{L_{t+1}}\delta\alpha k_{t}^{1-\alpha}}$ (C.3)
Collecting terms and using $\tau_{e}=\frac{A_{e}}{A_{m}}$, and noting that $\frac{N_{mt+1}A_{mt+1}}{L_{t+1}}$, the middle-aged share in total labor supply is constant over time and equal to $1-\gamma_{e}$, one less the elderly share, we have:
$\eta_{t+1}=\frac{\beta}{(1+\beta)\delta}-\frac{k_{t+1}}{k_{t}^{1-\alpha}}\left(\frac{\tau_{e}(1+a)}{(1-\alpha)(1+\beta)\delta}+\frac{(1+n)(1+a)}{\alpha\delta(1-\gamma_{e})}\right)$ (C.4)
This appears as (43) in the text. For equilibrium relationship (42), we begin by substituting the expressions in (40) into (38) to get:
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{A_{mt}k_{t}^{1-\alpha}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}+\frac{\frac{A_{et+1}k_{t+1}^{1-\alpha}}{R_{t+1}}}{\delta A_{mt-1}k_{t-1}^{1-\alpha}}\right)$ (C.5)
Using $A_{mt+1}=(1+a)A_{mt}$, $A_{mt-1}=\frac{A_{mt+1}}{(1+a)^{2}}$, and $\tau_{e}=\frac{A_{et}}{A_{m}t}$, gives:
$R_{t}=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}\left(\frac{k_{t}}{k_{t-1}}\right)^{1-\alpha}+\frac{r_{e}(1+a)^{2}}{R_{t+1}\delta}\left(\frac{k_{t+1}}{k_{t-1}}\right)^{1-\alpha}\right)$ (C.6)
Putting in (C.6) the expression for asset returns, $R_{t}=(1-\alpha)k_{t}^{-\alpha}$, gives expression (42) in the text. One convenience of equation (C.6) is that it yields a simple steady-state relationship for asset returns:
$R=\frac{1+n}{1+\beta}\left(\frac{1+a}{\delta}+\frac{(1+a)^{2}r_{e}}{\delta R}\right)$ (C.7)
which easily factors into equation (57) in the text.
Appendix D Proof of Proposition 7
In this appendix, we derive the relative consumption ratio for both the case in which inculcation occurs and when it does not, starting with the latter case. In this case, elderly consumption is current income plus income earned from asset accumulation in the prior period:
$c_{et}=y_{et}+s_{mt-1}R_{t}$ (D.1)
While for middle-aged agents, if there is no inculcative investment, current consumption is current income less savings:
$c_{mt}=y_{mt}-s_{mt}$ (D.2)
Moreover, one can use (37) to conclude that savings is determined by:
$s_{mt}=\frac{\beta y_{mt}-\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (D.3)
Plugging (D.3) into (D.2) gives:
$c_{mt}=\frac{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}{1+\beta}$ (D.4)
One can also use (D.3) to deduce that elderly consumptio in (D.1) can be written as:
$c_{et}=\frac{\beta}{1+\beta}\left(y_{et}+R_{t}y_{mt-1}\right)$ (D.5)
Forming the ratio of expressions and factoring out $R_{t}$ gives:
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\left(\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}\right)$ (D.6)
In the case where there is inculcation, we observe that consumption for the elderly is now:
$c_{et}=y_{et}+(1+n)g_{mt}+s_{mt-1}R_{t}$ (D.7)
while consumption for the middle-aged is:
$c_{mt}=y_{mt}-g_{mt}-s_{mt}-d_{t}\eta_{t+1}$ (D.8)
Elderly consumption is most easily found using (A.4), which allows us to conclude that:
$s_{mt-1}R_{t}+(1+n)g_{mt}+y_{et}=\beta R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (D.9)
Using (D.9) in (D.7) gives:
$c_{et}=R_{t}(y_{mt-1}-d_{t-1}\eta_{t}-s_{mt-1})$ (D.10)
Eliminating gifts in (D.8) using (36) gives:
$c_{mt}=(1-\eta)(y_{mt}-s_{mt}-d_{t}\eta_{t+1})$ (D.11)
For both (D.9) or (D.11), we can apply (37). Then, forming the relative consumption ratio gives:
$\frac{c_{et}}{c_{mt}}=\beta R_{t}\frac{y_{mt-1}+\frac{y_{et}}{R_{t}}}{y_{mt}+\frac{y_{et+1}}{R_{t+1}}}$ (D.12)
which is expression (44) in the text.